OpenStudy (anonymous):

find the slope and y-intercept of each line: 3x-4y=12

OpenStudy (anonymous):

have you seen the slop intercept form : y=mx+b?

OpenStudy (anonymous):

yes but how do i get i to y=mx=b form

OpenStudy (anonymous):

alright so first you want y by itself correct? what can you do to get y by itself

OpenStudy (anonymous):

and when i say y i mean -4y

OpenStudy (anonymous):

subtract the 3x

OpenStudy (anonymous):

yep now what do you get?

OpenStudy (anonymous):

you have -4y=12-3x -3 would be your slope and 12 is your y intercept?

OpenStudy (anonymous):

not exactly , the equation says y=mx+b, not -4y=mx+b. so you need one more step to get y's constant to be one

OpenStudy (anonymous):

divide by -4 ?

OpenStudy (anonymous):

yep =]

OpenStudy (anonymous):

so -3/4 and 3

OpenStudy (anonymous):

now you should have it in y=mx+b where m = slope and b is your y intercept

OpenStudy (anonymous):

12/-4=-3 -3/-4 = 3/4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

just a sign error but yes you have the right idea